-35-3x=-7(4x-7)-3x
Found 2 solutions by lwsshak3, stanbon:
Answer by lwsshak3(11628)

-35-3x=-7(4x-7)-3x
add 3x to both sides
-35=-7(4x-7)
divide both sides by -7
5=4x-7
4x=12
x=3

Answer by stanbon(75887)

-35-3x=-7(4x-7)-3x
-----
Add 3x to both sides to get:
-35 = -28x + 49
28x = 84
x = 3
--------------
Cheers,
Stan H.